Status Report

NOAA GOES SATOPS Morning Report 24 October 2000

By SpaceRef Editor
October 24, 2000
Filed under

SATOPS MORNING REPORT – October 24, 2000

* * Note:  the following report was revised to include a DMSP data
loss * *


No change in the status of the POES spacecraft.

NOAA-16 operations nominal.  The TED sweep voltage was turned off
for two revs this morning by request of the SEC.

This morning, the -X Instrument Mounting Platform (IMP) Thermal Control
Electronics (TCE) on NOAA-15 has been turned off in a continuing effort
to reduce the operating temperature of the AVHRR scan motor.  Yesterday,
the +X TCE was also turned off, with no appreciable change in the IMP thermal
environment.  Since the +X TCE was turned off, an increase in the
-X TCE duty cycle was observed, so it is hoped that the AVHRR scan motor
thermal environment will indeed decrease now that both the +/-X TCEs are
off.  The NOAA-15 AVHRR performance will be carefully evaluated over
the next 24-48 hours as the IMP temperature stabilizes.  Subsequent
to temperature stabilization, other options will be considered which may
include scan motor/electronics cycling and solar array movements.

NOAA-14 rev 29974 / F at 1648Z on October 23:  G5B received noisy
from spacecraft.  98% of the expected data received.  Data not
recoverable on subsequent revs due to recorder management.

POES operations were nominal over the past 24 hours.


No change in the status of the DMSP spacecraft.

F13/OPS47 rev 28818 (POGO/B at 07:41Z) had a 1360 spacecraft second
/ 34 playback second Category 1 data loss declared due to noisy data.

DMSP operations were nominal over the past 24 hours.


No change in the status of the GOES spacecraft.

GOES-8 images cancelled for KOZ – 0409, 0439, and 0509Z.  Last
night was the last night for KOZ operations for GOES-8.

The GOES-10 maneuver that was scheduled for Tuesday, JD-298 / 0ctober
24 was canceled due to space shuttle operations support requirements. 
It will be rescheduled for Friday, October 27 (maneuver time and maneuver
duration will be recalculated today).

GOES-8 East/West maneuver #32 is scheduled for October 25 (Day 299)
at approximately 1658Z.  The commanded thruster firing duration is
6.4 seconds.

Between 1115Z and 1846Z on Monday, October 23, multiple GOES-10 frame
breaks were observed and attributed to high BER at CDA.  As 1846Z,
GOES-10 support was shifted to the HR1 antenna.  Data overnight was
clean.  Actual frame breaks were recorded at 1115Z, 1119Z, 1121Z,
1131Z, 1135Z, 1136Z, 1140Z, 1145Z, 1201Z, 1230Z, 1245Z, 1252Z, 1300Z, 1315Z,
1321Z, 1330Z, 1345Z, 1352Z, 1400Z, 1415Z, 1422Z, 1430Z, 1445Z, 1452Z, 1500Z,
1600Z, 1735Z, 1741Z, 1830Z and 1846Z.

GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode
at ~105 deg W.

GOES operations were nominal over the past 24 hours.


Questions or comments to

SpaceRef staff editor.